OpenStudy (anonymous):

A desk is on sale for $150.80 , which is 71% less than the regular price. What is the regular price?

OpenStudy (anonymous):

(1 - 0.71)a = 150.80 solve for a 150.80/0.29 = 520

OpenStudy (xishem):

\[\frac{100-71}{100}=\frac{$150.80}{x}\]\[x=$520.00\]